OpenStudy (anonymous):

Solve: |x/3+1|=6 A.{-7, 15} B.{-21, 15} C.{7, 15} D.{-15, 21}

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

-21/3 = -7, and -7 +1 = -6, and |-6| = 6

OpenStudy (anonymous):

similarly, 15/3 = 5, 5+1 = 6, and |6| = 6

OpenStudy (anonymous):

\[|\frac{x}{3}+1|=6\] \[\frac{x}{3}+1=6\] or \[\frac{x}{3}+1=-6\] solve these

OpenStudy (anonymous):

b